OpenStudy (anonymous):

p^2-8p+7 still dont understand please explain!

OpenStudy (accessdenied):

What factors of 7 add up to get -8?

OpenStudy (anonymous):

Can you please post the full question. We are not clear what you are confused about.

OpenStudy (anonymous):

how to factor it

OpenStudy (anonymous):

Well one easy way to see it is if you have p^2 alone then you can factor it simply by (p) * (p) but we have all these extra things next to p^2 so we have to expand (p) * (p) so we can have it equal out

OpenStudy (anonymous):

@Daniela, In this case, 1 - 8 + 7 = 8 - 8 = 0 => 1 root p = 1, and other is 7 ( c/a = 7/1)

OpenStudy (anonymous):

There are methods to finding the extra stuff but I'm bad at explaining that so hopefully someone else can help

OpenStudy (anonymous):

@Romeo, you're much better explainer than I'm

OpenStudy (anonymous):

wait I got confused

OpenStudy (accessdenied):

I usually liked the "what factors of [constant term] add up to get [linear coefficient]?" explanation for when you have the leading coefficient of one on p*2 i.e. You have four factors of 7: 1, 7 and -1, -7 both work -7 and -1 add up to -8, which is our linear term. If we used these values in a factored form... (p - 7)(p - 1) p(p - 1) - 7(p - 1) then we can check through distributing again p^2 - p - 7p + 7 <--- those middle terms are our factors! p^2 - 8p + 7